# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems: A Computational Examination of the Korteweg de-Vries and Lorenz equations

### v1

@@ 4.1. Classical KdV equation
By Parseval’s identity $\norm{u}_{L^{2}}^{2}=2\sum_{k\in\mathbf{Z}}\absolutevalue{\hat{u}_{k}}^{2}=K^{2}$. Here $\epsilon$ is a specified constant, for computations we will use $\epsilon\sim e-16$ to denote machine precision. The purpose of finding such an $M$ is to separate the wavemodes into active low modes and inactive high modes. At any fixed time $t>0$ such an $M$ is guaranteed to exist as $u$ is analytic and therefore the Fourier coefficients must decay exponentially fast in wavenumber. In our computations we run only for a finite amount of time, so we take $M$ for $u$ to be the supremum over all such $M$’s taken across all simulated times. Without loss of generality we will assume that observation window consists of all $M$ modes, and so essentially the entire solution $u$ is known down to machine precision. We note here that $M$ is vital to the construction we outline below, thus examples we construct work only for finite times. This is due to the fact that we require a uniform bound on the number of active wavemodes to extend this arbitrarily far in time.
$\tilde{u}^{in}=c_{1}u^{in}+c_{2}\cos((2M+k)\pi x)$ (4.3)
Here $c_{1}$ and $c_{2}$ are positive constants chosen to enforce that $\norm{\tilde{u}^{in}}_{L^{2}}=K$ and $c_{2}>\epsilon$. That is, we are modifying the initial profile to take some of the momentum out of the active modes and shifting it into to wavemode $2M+k$, where $k$ is an arbitrary positive integer. As $k$ is arbitrary, we can construct infinitely many such solutions, all of which have the same momentum as the given solution, $u$, but with a new active frequency in the initial profile. Note that the choice of using ${u}^{in}$ is arbitrary, and one can do this exact construction using the solution of $u$ evaluated at any given time.
We performed computational tests using the reference solution, $\tilde{u}$, of Equation 2.6 with initial conditions prescribed by Equation 4.3. Here we took a solution, $u$, of Equation 2.6 which was initialized simply with $u^{in}=\cos(\pi x)$, which we found $M=50$ was suitable for the construction in Equation 4.3. Once $M=50$ was determined, we simply utilized $\tilde{u}^{in}=\cos(\pi\cdot x)+\cos(100\cdot\pi\cdot x)$, which was renormalized to ensure that $\norm{\tilde{u}^{in}}_{L^{2}}=\norm{u^{in}}_{L^{2}}$. We see in Figure 3 the result from applying nudging to attempt to recover the constructed reference solution $\tilde{u}$. Here the nudging solution, $v$, of Equation 2.7 is initialized to be identically zero. We see in Figure 3 that the low modes of $v$ synchronize with $\tilde{u}$ to a fixed level of precision, but the reference solution contains high frequency oscillations that $v$ never develops. The error in both the high and low modes appears approximately constant after a very brief initial decay in the observed modes. We note that in additional tests (not depicted here) the same behavior in the $L^{2}$ error was observed up to time $t=1000$ without any significant deviations in the error. One can clearly see activity in the frequencies immediately surrounding $k=100$ and $k=200$ that is simply not present in the nudged solution. Not only is this high frequency activity never generated, but nudging features no mechanism to generate targeted high mode oscillations. That is, if one knew a priori that the solution featured oscillations at wavemode $k=100$, but one still only had observations of the first $50$ wavemodes, nudging does not have any mechanism to encourage oscillations at any unobserved frequencies beyond the nonlinear term.
It is worth mentioning that the construction of the initial profile in Equation 4.3 is more complex than necessary. If one were to observe the first $M$ modes of a solution with a spike at wavemode $M+k$, $k$ being a positive integer, then nudging would likely fail. Here, when we say “spike”, we mean that looking at the $L^{2}$ spectrum one sees a local maximum at the specified wavemode. The construction of the initial profile in Equation 4.3 is done to produce examples where one can see by looking at the spectrum that nudging fails to develop any of the high oscillations featured in the reference solution. Indeed, we see in Figure 4 that we see similar behavior in the $L^{2}$ error to that seen in Figure 3. In Figure 4 we initialize the reference solution, $u$, simply with $u^{in}=\cos(\pi\cdot x)+0.001\cos(12\cdot\pi x)$, observing the first $10$ Fourier modes of $u$. The nudging solution, $v$, of Equation 2.7, was initialized to be identically zero. We see in Figure 4 that the nudged solution still fails to develop all the high frequencies present in the reference solution, even at time $t=1000$.
We note that the simulations presented here indicate that nudging is ill-suited to extremely capture high oscillatory solutions to Equation 2.6, but it does not necessarily prohibit other methods of data assimilation from recovering the reference solution. Unlike in Figure 2, where the observed data is identically zero, it is not impossible in theory for data assimilation to tell the difference between two solutions initialized of the form Equation 4.1. That is, unlike the case of Figure 2 if one initializes two solutions $u_{1}$ and $u_{2}$ of Equation 2.6 with initial data prescribed by Equation 4.1 with separate $k_{0}$ values for $u_{1}$ and $u_{2}$, we have that $\norm{P_{M}(u_{1}-u_{2})}_{L^{2}}\neq 0$ for $t>0$, and to the best of our knowledge one cannot readily construct two such solutions of this form that always agree on every observed wavemode for all time. This is due to the fact that dispersive term essentially causes the coefficients of the wavemodes to rotate in the complex plane around the origin with rotation speed $\delta^{2}k^{3}$. As each wavemode is rotating at a different speed one cannot rescale and shift momentum across wavemodes in a straightforward way such that the momentum backscattering due to the nonlinearity on the observed modes still agree at each timestep to produce a counterexample similar to the one present in Figures 2 and 1.
We note that all of these cases where nudging fails are specific to the undamped and unforced, classical KdV Equation 2.6, which is not a dissipative dynamical system. The addition of damping will result in the energy being drained from these high oscillations, while the forcing injects energy into specified low wavemodes, resulting in nontrivial long time dynamics that eliminate these sorts of extremely high frequency oscillations, see Section 4.2 below.
(a) Snapshot of solutions at time $t=10$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=10$.
(c) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(d) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 1. Nudging with $5$ observed modes initialized as in Equation 4.1 with $k_{0}=6$ fails to develop oscillations in the unobserved modes. Here $\mu=100$ and $\delta=0.075$.
(a) Snapshot of solutions at time $t=10$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=10$.
(c) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(d) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 2. Nudging with $5$ observed modes initialized as in Equation 4.1 with $k_{0}=6$ fails to control high oscillations towards the zero solution. Here $\mu=100$ and $\delta=0.075$.
(a) Snapshot of solutions at time $t=10$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=10$.
(c) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(d) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 3. Nudging with $50$ observed modes and additional activity at wavemode $k_{0}=100$, initialized as in Equation 4.3. Here $\mu=100$ and $\delta=0.075$.
(a) Snapshot of solutions at time $t=1000$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=1000$.
(c) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(d) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 4. Nudging with $10$ observed modes with $k_{0}=12$ fails to develop oscillations in the unobserved modes. Here $\mu=100$, $\delta=1$, and the reference solution is initialized as in Equation 4.3 with the $M$ chosen to be smaller than in Equation 4.2.
(a) Snapshot of solutions at time $t=1000$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=1000$.
(c) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(d) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 5. Nudging with $5$ observed modes with $k_{0}=3$ appears to work with an extremely slow rate of convergence. Here $\mu=100$ and $\delta=1$.
4.2. Damped and Driven KdV
$u_{t}+uu_{x}+\delta^{2}u_{xxx}+\gamma u=f,$ (4.4)
$v_{t}+vv_{x}+\delta^{2}v_{xxx}+\gamma v=f+\mu P_{M}(u-v).$ (4.5)
In [55] it was shown analytically, that given enough observed Fourier modes and large enough $\mu$, that the nudged solution converges to the reference solution in $L^{2}$ exponentially fast in time with exponent $\frac{\gamma}{4}$ for arbitrary choice of initial data $v^{in}\in H^{2}_{\text{per}}$ with spatial mean zero for Equation 4.5.
We now verify computationally that nudging successfully drives the assimilated solution to the reference solution exponentially fast in time. Depicted in Figure 6 we can see nudging effectively recovers a reference solutions initialized as in Equation 4.1 exponentially fast in time. Here we utilize $\mu=100$ and $\gamma=0.1$ and we note that the exponential rate of convergence is given by small perturbations from $\gamma$. Here we initialize the reference solution as in Equation 4.1 with $k_{0}=12$, that is, we selected the initial profile $u^{in}=\cos(12\cdot\pi\cdot x)$. For our driving force, we utilized $f=\cos(8\cdot\pi\cdot x)$. We note that and we observed only the first 10 modes of the solution. It is unsurprising that convergence happens at the exponential rate proportional to the damping coefficient, $\gamma$, as this coincides with the results of [55]. We note that the exponential convergence is not entirely independent of $\mu$, the oscillatory component at initial times and on the observed modes is particularly sensitive to $\mu$ values, although this effect quickly fades as the error begins to decay at rate $\gamma$. Moreover, as one sees from the spectrum plots in Figure 6 the primary barrier for convergence are the spikes in undetected frequencies, therefore it is unsurprising that convergence to the reference solution happens at exactly the rate, $\gamma$, at which momentum is drained from these modes. Indeed, this behavior is typical of simulations of nudging methods, as the unobserved modes rely on the dissipation in the system in order to synchronize. The overall decay in error can typically be decomposed into two distinct phases, an initial rapid decay in error that is highly dependent on $\mu$, and a slower, but still exponentially fast in time, rate of convergence primarily driven by the dissipation in the system. For an in-depth study of the role of the particular value of $\mu$ see, e.g., [15].
We note that the AOT algorithm successfully recovers the reference solution, $u$, to Equation 4.4 even when $u$ is initialized as in the counterexamples from the previous section Equations 4.1 and 4.3 given enough observed modes. We also point out, that in Figure 6 we see some behavior in the observed modes due to the presence of the body force. We could disable the body force by setting it to 00, however in that case data assimilation would not be required at all. This is because the damping term will drive any initial condition to 00, exponentially fast at the rate $\gamma$ with or without any form of data assimilation being applied.
(a) Snapshot of solutions at time $t=10$.
(b) Snapshot of $L^{2}$ spectrum of solutions at time $t=10$.
(c) Snapshot of solutions at time $t=300$.
(d) Snapshot of $L^{2}$ spectrum of solutions at time $t=300$.
(e) $\norm{u-v}_{L^{2}}$ over time split into observed low modes, unobserved high modes, and the total error.
(f) Approximate exponential decay rate of the high mode error for the last $500$ timesteps.
Figure 6. Nudging with $10$ observed modes successfully recovers reference solution with nudging with $\mu=100$ for $\gamma=0.1$, $\delta=1$, and $f=\cos(8\pi x)$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation 4.1 with $k_{0}=12$. Note that the solution is not converging to zero, but features small amplitude oscillations.
5. Lorenz 1963 System
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{Z}=-bZ+XY-b(r+\sigma),&Z(0)=Z_{0}.\end{cases}$ (5.1)
It is also known that applying the AOT algorithm, or a synchronization scheme, as in [49, 48], by observing only the $X$ component is sufficient to synchronize the nudged solution with the reference solution (see e.g. [49, 12, 48]). In particular, we can consider the following three synchronization schemes from [48] (which was motivated by [68]) to understand when data assimilation works for this system.
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{y}=-\sigma X-y-Xz,&y(0)=y_{0}\\ \dot{z}=-bz+Xy-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (5.2)
$\begin{cases}\dot{x}=-\sigma x+\sigma Y,&x(0)=x_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{z}=-bz+xY-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (5.3)
$\begin{cases}\dot{x}=-\sigma x+\sigma y,&x(0)=x_{0}\\ \dot{y}=-\sigma x-y-xZ,&y(0)=y_{0}\\ \dot{Z}=-bZ+XY-b(r+\sigma),&Z(0)=Z_{0}.\end{cases}$ (5.4)
The above equations, Equations 5.2, 5.3 and 5.4, correspond to the regimes when the observed values of $X$, $Y$, or $Z$, respectively, are observed and inserted directly into the equations for the remaining, unobserved components. That is, e.g. in Equation 5.2 we observe the true value of $X$ continuously in time and insert that into the evolution equations of $y$ and $z$. In these equations we note that numerically we should not be simulating the evolution of $X$, $Y$, or $Z$ as they appear in Equations 5.2, 5.3 and 5.4, respectively, but rather these values are observed (given) from the true solution in Equation 5.1 and are inserted directly into the equations for the remaining variables. Analysis on Equations 5.2, 5.3 and 5.4 was conducted in [48] where it was shown that the continuous coupling on $X$ or $Y$ alone was sufficient to obtain synchronization between the assimilated equations and the reference solution, however the coupling seen in Equation 5.4 was not sufficient to obtain convergence in either the $x$ or $y$ components to the corresponding components $X$ or $Y$ of the true solution. We do note that in [48] the equations for Equation 5.1 were written in a slightly different form, however this can be easily reconciled by using the transformation $\tilde{Z}=Z+(r+\sigma)$.
For the sake of completeness, we will discuss the convergence results for the above systems, Equations 5.2, 5.3 and 5.4 proven in [48] and discuss how the results break down in the absence of dissipation. Let
$\delta=(X-x,Y-y,Z-z),$ (5.5)
where $(X,Y,Z)$ is the given solution of Equation 5.1 and $(x,y,z)$ is the solution to one of the assimilated equation, Equations 5.2, 5.3 and 5.4. Here we define $\delta$ with the understanding that the variable in $(x,y,z)$ missing from the assimilated equation is taken to be equal to the corresponding variable in Equation 5.1, e.g. when considering Equation 5.2 we define $x(t):=X(t)$ for all time, $t$ and have $\delta=(0,Y-y,Z-z)$. The convergence results for the synchronization algorithms, Equations 5.2, 5.3 and 5.4 are given as follows:
Theorem 5.0.1 (Hayden (2007) [48]).
Let $(X,Y,Z)$, be the solution to Equation 5.1, then the following are true when $\sigma>0,b>0,$ and $r>0$.
If $(x,y,z)$ is the solution to Equation 5.2, then
$\norm{\delta}^{2}_{\ell^{2}}\leq\norm{\delta_{0}}_{\ell^{2}}e^{-2t}.$ (5.6)
If $(x,y,z)$ is the solution to Equation 5.3, then
$\norm{\delta}^{2}_{\ell^{2}}\leq|X(0)-x(0)|e^{-2\sigma}+|Z(0)-z(0)|e^{-2b}+\frac{J|X(0)-x(0)|^{2}}{b}\left(e^{-bt}-e^{-2\sigma t}\right),$ (5.7)
where
$J=\begin{cases}r^{2},&b\leq 2\\ \frac{b^{2}r^{2}}{4(b-1)},&b\geq 2\end{cases}$ (5.8)
The proofs of 5.0.1 are derived by taking the difference of the assimilated equation (Equation 5.2 or Equation 5.3) with Equation 5.1, summing the squares of the derivatives, and applying Grönwall’s inequality to the resulting quantity. The $J$ estimate in Equation 5.7 arises from the use of the global bound
$Y^{2}\leq J,$ (5.9)
which was derived in [28] and was used to bound the contribution from the nonlinear term in the $z$ equation.
We note that the analysis done in the proof of 5.0.1 is dependent on the particular parameters and requires $\sigma>0$, $b>0$ and $r>0$. To illustrate how these sorts of estimates break down in the absence of dissipation, we consider the case when both $X$ and $Y$ are observed, i.e., are given from the true solution of Equation 5.1. This leads to the following system of equations
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{z}=-bz+XY-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (5.10)
Now, if we take $w=z-Z$, where $Z$ solves Equation 5.1, we see from Equations 5.1 and 5.10 that $w$ satisfies the following equation
$\dot{w}=-bw,\quad w(0)=z_{0}-Z_{0}.$ (5.11)
$|w(t)|^{2}\leq|w(0)|^{2}e^{-bt},$ (5.12)
which yields convergence of $z$ to $Z$ exponentially fast in time. If we instead consider $b=0$, removing the dissipation on the $Z$ component of Equation 5.1, then we now have that $\dot{w}=0$, and so $w(t)=z_{0}-Z_{0}$ holds for all time. Therefore, we have that the absence of dissipation on the $Z$ component prevents synchronization between the assimilated solution and the reference solution, even when both the $X$ and $Y$ components are observed continuously in time.
We note that the above discussion has been on the synchronization data assimilation scheme, given by the direct insertion, as it was first introduced in [68], of the observed quantities into the model equations. We now consider the nudged Lorenz system for the observed $X$ and $Y$ components of Equation 5.1, given as follows:
$\begin{cases}\dot{x}=-\sigma x+\sigma y+\mu(X-x)\\ \dot{y}=-\sigma x-y-xz+\mu(Y-y)\\ \dot{z}=-bz+xy-b(r+\sigma)\end{cases}$ (5.13)
At any fixed time, $t$, the solutions, $(x(t;\mu),y(t;\mu),z(t;\mu))$, to the nudged equations, Equation 5.13, are expected to converge to the solution $(X(t),Y(t),z(t))$ of the synchronization algorithm, given by Equation 5.10 as $\mu\to\infty$. This can be justified rigorously by applying the argument present in [15] to the Lorenz 1963 system, provided both methods are initialized with the same initial data. An analogous statement was made in [72] for the Lorenz system in particular (with $b>0$), where it was shown that the solution, $x(t;\mu)$, of Equation 5.13, converges to $X(t)$ at any fixed time, $t$, as we send $\mu\to\infty$ when the nudging term is present only on the equation for $x$ in Equation 5.13. In the case where $b=0$, if we assume a priori that the solutions of Equation 5.13 converge to those of Equation 5.10 as $\mu\to\infty$, one can deduce using the reverse triangle inequality, that $\norm{\delta(t)}^{2}_{\ell^{2}}\geq|z_{0}-Z_{0}|^{2}$, where $\delta$ is given as defined previously in Equation 5.5. Therefore, we expect the nudged Lorenz system, Equation 5.13, to fail to recover the true solution in the same way that Equation 5.10 fails when dissipation is removed from the equations for $z$.
We now demonstrate numerically that this dissipation mechanism (when $\sigma,b>0$) of the Lorenz system is necessary for data assimilation methods to work. In particular, we will be considering a partially dissipative variant of the Lorenz system by taking $b=0$, i.e., eliminating the dissipation mechanism only on the $Z$ component. The analytical results regarding the nudging algorithm applied to the Lorenz equations have been done when $b>0$ and are qualitatively similar to those we reference above in 5.0.1 from [48], see e.g. [12, 72, 17]. That is, the solutions of Equation 5.13 synchronize with those of Equation 5.1 exponentially fast in time for large enough nudging parameter values and with the exponential rate depending on the particular parameters of Equation 5.1. While nudging is only required on either the $X$ or $Y$ components in order to drive convergence, we will utilize nudging on both the $X$ and $Y$ components to emphasize that, even with additional nudging, the method is not effective when the dissipation on the $z$-component is eliminated (i.e. when $b=0$).
First, we verify the validity of nudging on the Lorenz system with full dissipation. Using the initial data and parameter choices discussed previously we apply nudging to the Lorenz system with $b=8/3$. The results can be seen in Figure 7, where we have applied nudging to only the first two components of Equation 5.13. Here we can see that all variables converge to an error level of approximately $121e-$ 12 with spurious dips below this error level in each component. Note that here we see exponential convergence to the true solution, as we expect in this case.
Next, we examine what happens when we apply nudging for the partially dissipative Lorenz system, i.e., with $b=0$. The results can be seen in Figure 8. Note that here we still see exponential decay to the true solution in both $x$ and $y$. While nudging appears to affect $z$ initially, this effect quickly fades after a short interval of time and the error appears to stabilize to be approximately constant $11e+$ 1. The error level that is attained appears to be directly linked to the difference in the initial conditions for the $z$ component. To illustrate this, we plotted the $\ell^{2}$ (Euclidean norm) error over time when both the reference and nudged equations are initialized with the same initial data, except for the $z$ component. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, where $\delta$ ranges in values from $11e-$ 1 to $121e-$ 12. The results of these tests can be seen in Figure 9 and they illustrate that the $\ell^{2}$ error is highly sensitive to small perturbations in the initial condition for $z$. We note that the behavior seen in Figure 9 appears to align with what one should expect from our earlier discussion of Equation 5.11 for the synchronization algorithm. We note that we obtain an initial exponential decay in the error in the $z$ component which is driven by the convergence in both $x$ and $y$. This initial convergence levels off and becomes approximately constant as $x$ and $y$ converge to $X$ and $Y$, respectively.

### v2

@@ 4.1. Analytical Results
$\begin{split}\boldsymbol{v}_{t}+B(\boldsymbol{v},\boldsymbol{v})&=\mu P_{\sigma}P_{M}(\boldsymbol{u}-\boldsymbol{v}).\end{split}$ (37)
Here we formally denote the nonlinear operator, $B(\boldsymbol{a},\boldsymbol{b})\mathrel{\mathop{\ordinarycolon}}=P_{\sigma}(\boldsymbol{a}\cdot\gradient\boldsymbol{b})$.
We recall the following theorem regarding global well-posedness of solutions to Section˜4.1
Theorem 4.1 (Global well-posedness of Section˜4.1 (see e.g. [64])).
Given $\boldsymbol{u}^{in}\in V^{s}(\mathbb{T}^{2})$ with $s\geq 3$, then Section˜4.1 is globally well-posed with solution
$\boldsymbol{u}\in C([0,T];V^{s})\cap L^{\infty}([0,T];H).$
Moreover $\mathinner{\!\left\lVert\boldsymbol{u}(t)\right\rVert}_{H}=\mathinner{\!\left\lVert\boldsymbol{u}^{in}\right\rVert}_{H}$ for all $t\geq 0$.
We now move to restate our propositions from Section˜3.2.1 in this context.
Proposition 4.2.
Let $k\in\mathbb{Z}\setminus\{0\}$ and $\phi\in L^{2}_{\text{per}}$ with $\phi$ having spatial mean zero. The following hold true:
  1. (i)
    $\phi(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e_{i}})=\phi(\boldsymbol{x})$ holds for a.e. $x\in\mathbb{R}^{2}$ for $i=1,2$ if and only if
    $\phi(\boldsymbol{x})=\sum_{\boldsymbol{n}\in k\mathbb{Z}^{2}}\hat{\phi}_{\boldsymbol{n}}e^{i\frac{2\pi}{L}\boldsymbol{n}\cdot\boldsymbol{x}}.$ (38)
  2. (ii)
    Let $M$ be a positive integer and let $k$ be an integer satisfying $\mathinner{\!\left\lvert k\right\rvert}\geq M$. Suppose $\phi(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\phi(\boldsymbol{x})$ for a.e. $x\in\mathbb{R}^{2}$ for $i=1,2$, then
    $P_{M}\phi=0$ (39)
Proposition 4.3.
Suppose $k\in\mathbb{Z}\setminus\{0\}$ and $\boldsymbol{u}^{in}\in V^{3}$ such that $\boldsymbol{u}^{in}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\boldsymbol{u}^{in}(x)$ for all $x\in\mathbb{R}$ and each $i=1,2$. Additionally suppose that $\boldsymbol{u}^{in}$ has spatial mean zero. Let $u(\boldsymbol{x},t)$ be the corresponding solution of Section˜4.1, subject to periodic boundary conditions with period $L$ in both the horizontal and vertical directions.
Then $u(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=u(x,t)$ for a.e. $x\in\mathbb{R}$ and all $t\in\mathbb{R}$.
Proof.
Let $\boldsymbol{v}(\boldsymbol{x},t)=u(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)$ for a fixed $i=1$ or $2$ and notice that
$\boldsymbol{v}(\boldsymbol{x},0)=\boldsymbol{v}^{in}(\boldsymbol{x})=\boldsymbol{u}^{in}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},0)=\boldsymbol{u}^{in}(\boldsymbol{x}).$
Thus we see that $\boldsymbol{v}(\boldsymbol{x},t)$ is also a solution of Section˜4.1. As $\boldsymbol{v}(\boldsymbol{x},0)=\boldsymbol{u}^{in}(\boldsymbol{x})$, then by the uniqueness of solutions to Section˜4.1 we obtain that $\boldsymbol{v}(\boldsymbol{x},t)=\boldsymbol{u}(\boldsymbol{x},t)$, hence we have that $\boldsymbol{u}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=\boldsymbol{u}(\boldsymbol{x},t)$. ∎
Corollary 4.4.
Under the conditions of Proposition˜3.3 the space of functions periodic with respect to the lattice $k\mathbb{Z}^{2}$ is invariant under the solution of Section˜4.1.
Proposition 4.5.
Suppose $k\in\mathbb{Z}\setminus\{0\}$ is given and suppose $\boldsymbol{u}^{in}_{j}\in V^{3}$, for $j=1,2$, satisfying
$\boldsymbol{u}^{in}_{j}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\boldsymbol{u}_{j}^{in}(\boldsymbol{x})$ (40)
for a.e. $\boldsymbol{x}\in\mathbb{R}^{2}$, $i=1,2$, and $\boldsymbol{u}_{j}^{in}$ has spatial mean zero. Moreover, assume $\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{1}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{2}\right\rVert}_{L^{2}}$ and let $\boldsymbol{u}_{j}(x,t)$ be the corresponding solutions of Section˜4.1. Then
  1. (i)
    $\boldsymbol{u}_{j}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=\boldsymbol{u}_{j}(\boldsymbol{x},t)$ for a.e. $\boldsymbol{x}\in\mathbb{R}$, $i=1,2$, and for $j=1,2$ and all $t\in\mathbb{R}^{2}$.
  2. (ii)
    $P_{M}\boldsymbol{u}_{1}(\cdot,t)=P_{M}\boldsymbol{u}_{2}(\cdot,t)$ for any positive integer $M<\mathinner{\!\left\lvert k\right\rvert}$, and every $t\in\mathbb{R}^{2}$.
  3. (iii)
    $\limsup\limits_{t\to\infty}\mathinner{\!\left\lVert\boldsymbol{u}_{1}(\cdot,t)-\boldsymbol{u}_{2}(\cdot,t)\right\rVert}_{L^{2}}>0.$
Now, using these propositions, we show that Section˜4.1 does not have the finitely many determining modes property.
Theorem 4.6.
The incompressible Euler equations Section˜4.1 subject to periodic boundary conditions on $\mathbb{T}^{2}$ does not possess the finitely many determining modes property in $H$.
Corollary 4.7.
For a fixed number of observed Fourier modes, $M$, there exists solutions $\boldsymbol{u}(\boldsymbol{x},t)$ of Section˜4.1 that cannot be recovered by Section˜4.1 with arbitrary initial condition $\boldsymbol{v}^{in}$.
Proof.
Let $M\in\mathbb{N}$ be given and let $\boldsymbol{u}(\boldsymbol{x},t)$ be any nontrivial solution to Section˜4.1 that is periodic with respect to the lattice $k\mathbb{Z}^{2}$ for $k\in\mathbb{Z}$ given such that $\mathinner{\!\left\lvert k\right\rvert}>M$.
Now, as $\boldsymbol{v}^{in}$ is arbitrary we can specify $\boldsymbol{v}^{in}\neq\boldsymbol{u}^{in}$ with $\boldsymbol{v}^{in}$ also being periodic with respect to the same lattice as $\boldsymbol{u}$. We will additionally assume that $\mathinner{\!\left\lVert\boldsymbol{u}^{in}\right\rVert}_{H}\neq\mathinner{\!\left\lVert\boldsymbol{v}^{in}\right\rVert}_{H}$.
We note that, by construction, the nudging term $\mu P_{\sigma}P_{M}(\boldsymbol{u}-\boldsymbol{v})$ drops out entirely in Equation˜37, and thus one can easily verify that $\boldsymbol{v}(\boldsymbol{x},t)$ is in fact a solution to Section˜4.1 corresponding to the initial data $\boldsymbol{v}^{in}$. The result immediately follows from the fact that the $H$ norm is a conserved quantity for solutions to Section˜4.1. ∎
4.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Sections˜4.1 and 4.1. All simulations depicted in this study were performed Matlab (R2023a) using our own code. For this computational study we implemented the incompressible Euler equations on the 2D periodic torus $\mathbb{T}^{2}=[0,2\pi)\times[0,2\pi)$, subject to periodic B.C. with period with pseudo-spectral methods. We implement the equations by using the explicit Euler time-stepping scheme with the nonlinear computed explicitly. In the simulations in this chapter we utilize pseudo-spectral methods (see e.g., [14, 72, 75] for textbook descriptions of these methods), that is, derivatives were computed using Matlab’s implementation of the n-dimensional fast Fourier transform (fftn, which relies on FFTW). We utilized the standard 2/3 dealiasing rule to compute the nonlinear term, that is, we zeroed out the modes above wavenumber $N/3$, where $N$ is number of gridpoints in one direction of the periodic box. Moreover, to reduce the number of FFT operations, the Basdevant formula was used (see [30] for further discussion). We utilize the spatial resolution $N=2^{8}$ and time step $\Delta t=0.001$.
We remark that while our analytical results illustrate the failure of nudging to recover the true solution, effective testing of the AOT algorithm in recovering solutions to Section˜4.1 remains problematic. This is due to the fact that the AOT algorithm is typically shown to recover dissipative dynamical systems only asymptotically in time. Thus if one wishes to faithfully demonstrate the effectiveness of the AOT algorithm it requires running simulations of Section˜4.1 for long amounts of time. This in itself is problematic, as it was shown in [29] that solutions to Section˜4.1 supported on a finite number of Fourier modes for all times $t\in[0,\infty)$ must necessarily be constant in time. This indicates that solutions that are evolving in time develop arbitrary fine length scales and thus one cannot accurately simulate the long time dynamics using a fixed spatial resolution. However, we note that steady-state solutions are in fact enough to illustrate the failing of the AOT algorithm on the incompressible Euler equations.
For our reference solution we utilize the Taylor-Green vortex given with frequency $k=15$ and $c=10^{-4}$, given as follows
$\begin{cases}u_{1}(\boldsymbol{x},t)=c\sin(k\boldsymbol{x}_{1})\cos(k\boldsymbol{x}_{2}),\\ u_{2}(\boldsymbol{x},t)=-c\cos(k\boldsymbol{x}_{1})\sin(k\boldsymbol{x}_{2}).\end{cases}$ (41)
We initialize the nudged equation with initial data identically zero $\boldsymbol{v}^{in}\equiv\boldsymbol{0}$, with $\mu=100$. The results of our simulations can be seen below in Figures˜11 and 12.
We see in Figure˜11(a) that observing all Fourier modes $\boldsymbol{k}$ with $\mathinner{\!\left\lvert\boldsymbol{k}\right\rvert}\leq 21$ is insufficient to recover the true solution. We note that in the spectral plots Figures˜11(b) and 12(b) the vertical lines indicate the observational frequency and dealiasing cutoff frequencies, given in blue and red, respectively. While $M=21$ is insufficient to recover the true solution, we find that $M=22$ is sufficient, which is of course unsurprising as this is exactly the regime in which the entire solution is in fact observed. When $M$ is chosen smaller than the frequency utilized in the construction of the Taylor-Green vortex the solution $\boldsymbol{v}(\boldsymbol{x},t)\equiv\boldsymbol{0}$ holds for all time $t\in[0,1]$.
Figure 10. Reference solution given by Taylor-Green steady state with active frequency $k=15$ and amplitude $c=10^{-4}$.
((a)) Error over time for $\mu=100$ with $M=21$ observed modes.
((b)) Snapshot of $L^{2}$ spectrum of solutions at time $t=1$.
Figure 11. Nudging with 2121 observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜41 with $k=15$. Note that $M=21$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$
((a)) Error over time for $\mu=100$ with $M=22$ observed modes.
((b)) Snapshot of $L^{2}$ spectrum of solutions at time $t=1$.
Figure 12. Nudging with $22$ observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜41 with $k=15$. Note that $M=22$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$